# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v1

@@ 2.3.3 Our contribution: Is the language model using the elicited prior in-context?
To quantify this, we use the energy statistic [33] (Appendix A.4), which measures how closely two distributions match (taking values $\in[0,1]$ with 00 indicating two distributions are equal). Here, we test the difference between the parameter distributions from either the elicited distribution ($\theta\sim p_{M,T}(\theta)$) or the one approximated from the language model’s in-context predictive model ($\varphi\sim p_{M,T}(\varphi)$).
@@ 2.3.4 Our contribution: Inspecting the in-context posterior distribution
If the language model’s internal posterior is significantly different to the one calculated by applying Monte Carlo methods to the approximated prior, then in-context learning might be considered unusable for a given task, as it is not clear that the model is updating its prior belief about a predictive task from the data. This would suggest that the language model should not be used to generate posterior distributions on predictions in-context as it is improperly applying Bayes rule after observing demonstrations.
@@ 2.4 Bayes factor for model selection
To decide, given a dataset to use in-context learning or prior elicitation, we consider the Bayes factor for model selection.
The Bayes factor compares the ratio of marginal likelihoods of two statistical models to determine under which model the data was most likely generated. Formally, given two models $\alpha_{0}$ and $\alpha_{1}$ with parameterisations $\theta_{0}$ and $\theta_{1}$, the Bayes factor calculated over a dataset $\mathcal{D}$ is:
@@
In our case, $\alpha_{0}$ refers to prior elicitation, with the prior determining the parameterisation $\theta$ of a predictive model, and $\alpha_{1}$ referring to in-context learning. We can approximate the Bayes factor by using the means from samples of the likelihood of $D$ (Section 3.6).
We chose the Bayes factor over the Bayesian information criterion (BIC) because BIC uses the number of model parameters as part of the model selection. When comparing prior elicitation with in-context learning, BIC would almost always favour prior elicitation because the number of parameters required to make a prediction using in-context learning is on the order $10^{9}$.
@@ 2.5 Experimental datasets
Within our experiments we present results on five real-world datasets and a synthetic dataset, covering both classification and regression tasks from various domains.
@@
To demonstrate the effectiveness of the LLM elicited priors on new datasets, we use a private (i.e., not seen by the LLM) dataset collected as part of a study on Dementia care which contains $10$ features of in-home activity and physiology and labels corresponding to days with a clinically validated labelling of positive or negative urinary tract infection (UTI), calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ positive or negative days of UTI. This dataset has not appeared online and will not have been part of the language model’s training.
We then test four public benchmark datasets. The Breast Cancer dataset [36], where the task is to predict the diagnosis of a breast mass; the California Housing dataset [37], with the goal to predict the median house value from the U.S census data for that region; the Wine Quality dataset [38, 39] where the task is to predict the wine quality from its physicochemical characteristics; and the Heart Disease dataset [40, 41] where the task is to predict a heart disease diagnosis from physiological features.
@@ 3 Results
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo, with further language models tested in Appendix A.13.1 and A.13.2.
@@ 3.1 Evaluating the LLM-elicited priors
On the synthetic data, we learn to model $y=2x_{1}-x_{2}+x_{3}$ using a prior elicited from the language model. In this case, as we are testing how well the language model can provide priors based on informative task descriptions, we describe the exact relationships between $x$ and $y$ in the task descriptions. The LLM-elicited prior provides a lower MSE, by an order of magnitude, compared to an uninformative prior on the test set after $5$ observed examples and an equivalent MSE after $20$ data points are seen. This is likely because the task was simple enough to be learnt effectively by an uninformative prior after a few examples. However, it demonstrates that the language model is able to provide effective prior distributions given a clear task description. In fact, the mixture of the priors that the language suggests are shown in Appendix A.6.1 and are centred on the true parameter values, suggesting that the language model is able to convert its combined knowledge of a problem and a task description into useful prior distributions. The relationship between the detail of descriptions and the performance of elicited priors is studied within Appendix A.10.
For the real world datasets, we often see the same posterior performance improvements. On the Breast Cancer dataset, the improvement in accuracy is clear up to $40$ observed data points, where the mean accuracy is greater, and its variation is smaller. On the other hand, for the Wine Quality dataset, Figure 2(b) shows that the prior distribution on $\theta$ elicited from the language model is too strict as it eventually reduces the mean accuracy of the logistic regression model when the number of training points is larger than $25$, whilst the variation in accuracy remains smaller. Given these samples, the likelihood of the prior distributions could have been measured (as is described in 2.4) before the elicited distribution was used for new predictions.
@@
For both the California Housing and UTI datasets, the elicited prior provides a greater mean accuracy over the uninformative prior for all of the number of observed data points tested, illustrating the potential of using a language model for prior elicitation. This should be of particular note considering that the UTI dataset is private, and will have never appeared within the language model’s training.
In general, we see that the greatest improvement in accuracy occurs for smaller numbers of observed samples, suggesting that using prior knowledge from language models is most relevant when users have access to only a few samples.
This can be important in scenarios when collecting labels is costly and time-intensive, as is the case in the UTI dataset. In Appendix A.6.2, we consider the accuracy achieved using uninformative priors and elicited priors in the context of collection time. We show here that a UTI predictive model might have achieved greater accuracy earlier in the study. In a clinical environment, this could have enabled improved UTI detection earlier, accelerating the model development and validation.
In Appendix A.6, we present further accuracy distributions (similar to Figure 2(a)), using the Heart Disease and Wine Quality datasets, which supplement the results in Figure 2(b). We further test the number of task descriptions required in Appendix A.11, and the quality of those descriptions in A.10.
Given two sets of priors, one elicited from a language model and an uninformative prior, the prior predictive distribution can be sampled and the likelihoods calculated given some data points. This can be done similarly to that described in Section 3.6.
@@ 3.2 Incorporating expert information
To study this, we chose to additionally include our own knowledge of the risk factors for UTI in the task description for prior elicitation [43]. In particular, we provided further information about three features available in the dataset which are positively correlated with outcome. We added the following phrasing to the system prompts, based on clinical knowledge of UTI: “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
@@
It is interesting however, that without this expert information, the LLM had already correctly identified the relationship between UTI risk and the bathroom frequency, night-time disturbances, and the number of previous UTIs, as these are all positive predictors of UTI. The language model had already translated this information from its internal prior into the elicited distributions. This could explain the small improvement in the posterior accuracy over the LLM-elicited priors without additional expert information, shown in the line graph of Figure 3.
@@ 3.3 Extracting the in-context prior and posterior
The prior and posterior distributions estimated through a repeated sampling of the in-context predictions (as outlined in Section 2.3.2) are shown in Figure 4 and 5. Once we estimate the parameters of the linear model, we can use them to calculate the logits (or regression values) of the predictions returned by the language model (from in-context learning) and measure the error between the outputs of the language model and the estimated internal model. This is shown in the bottom row of Figure 4 and 5, which suggests that our estimations of the prior internal parameters are reliable for the real-world datasets. The upper row presents the distributions of the estimated parameters, with each colour representing a single feature.
For the prior distributions in Figure 4, we see variations in the estimated prior distributions across the datasets, suggesting that the language model is using some knowledge about tasks when making predictions, even without any demonstration examples. Even so, we see that the language model is often using a prior centred around 00, suggesting it is not particularly informative.
The extracted posterior distributions are given in Figure 5 and show some differences compared to the prior distributions. For California Housing, Heart Disease, and Wine Quality, our estimation error increased, but remained low. However, we find that the Breast Cancer dataset produces a large error for our predictions of the logits (Figure 5), suggesting that our estimates of the parameters of the internal model are not reliable for this case. It is interesting that on the whole, we were able to produce a better approximation of the language model’s internal prior distribution than the posterior distribution, and is likely due to the limitations of the language model and in particular, its capacity to understand demonstrations and instructions in the same prompt.
@@ 3.4 Is the language model being truthful?
Table 1 illustrates the differences between the extracted in-context prior distribution (Section 2.3.2) and the LLM-elicited distribution (Section 2.2). No dataset has similar distributions according to their energy distance, demonstrating that the language model is using a different prior distribution to make its own predictions compared to the one it provides for us to construct a separate predictive model.
These conclusions are supported in Appendix A.7, where we present histograms of the parameter values for each feature, which show that the two distributions significantly differ from each other visually, with few similarities. This suggests that either the language model is unable to provide us with a reliable representation of its own prior knowledge or that it is unable to reliably use its own prior knowledge when making predictions. Since the in-context prior distributions are often centred around 00, we suspect that it is the latter, and that the language model is not effectively using its own prior beliefs about a task in-context.
Since these priors disagree, we will investigate them further in Section 3.5, where we explore empirically whether in-context learning is approximating Bayesian inference and in Section 3.6, where we study decision criteria for understanding which of these distributions is more likely to have generated the observed data using the Bayes factor.
@@ 3.5 Is in-context learning approximating Bayesian inference?
It is important that the model is performing Bayesian inference, so that newly observed data points and the prior knowledge are combined to form a posterior predictive distribution. In this section, we empirically evaluate the extent to which the language model is performing Bayesian inference in-context, by measuring the difference between the extracted in-context posterior distribution and Monte Carlo sampling applied to the extracted prior in-context distribution (Section 2.3.4).
@@
On the other hand, for Wine Quality and Heart Disease, the difference between the extracted posterior and the posterior samples of the extracted prior match closely, and the error in our estimations of the parameters are low (Figures 4 and 5). In these cases, we are more confident that in-context learning is imitating Bayesian inference and so should use a model selection tool to decide between prior elicitation and in-context learning.
Appendix A.8 presents histograms of each feature’s parameter values after Monte Carlo sampling or in-context learning have been performed. Figure 12 within Appendix A.8 supports the findings in this section and allows us to visualise the differences illustrated by Table 2.
These results provide empirical evidence that in-context learning is sometimes approximating Bayesian inference but not necessarily for all tasks.
@@ 3.6 Bayes factor and model selection
For in-context learning, we do not have access to $\varphi$ of the internal predictive model. However, since the language model provides us with predictions on targets given features directly, we can still approximate $p(\mathcal{D}|\alpha_{1})$ through a repeated sampling of predictions. Here, we prompt the language model with our $100$ task descriptions and ask it to predict on $25$ data points from our dataset. We repeat this with $5$ different splits of the dataset to calculate the mean and standard deviation of the Bayes factor.
@@
For the regression tasks, prior elicitation is preferred for $y=2x_{1}-x_{2}+x_{3}$, whilst for California Housing in-context learning produces greater log-likelihoods. However, here, both in-context learning and prior elicitation provide a high prior mean squared error and low marginal likelihood, making the Bayes factor unreliable to compute. Further, in the case of California Housing in-context learning is not reliably estimating Bayesian inference (Section 3.5), suggesting prior elicitation may still be preferable.
Whether prior elicitation or in-context learning provides prior distributions with greater likelihoods depends on the task, with both methods being the better choice over a non-informative prior for all cases tested. However, when factoring in the methods’ abilities to estimate Bayesian inference and the cost of producing predictions, prior elicitation is likely the better option for the cases tested.
@@ 3.7 Choosing between priors
As seen in Appendix A.13.1 and A.13.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior, however this might not be the case for all datasets or language models. Given prior distributions elicited from different language models, we would like to choose the best one for our predictive task. For this, given some collected data, we can use the Bayes factor (Section 2.4) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques such as the Bayesian information criterion. This is because we are performing the Bayesian inference ourselves, given the prior distributions. Further, this method also allows us to compare the elicited distributions to the uninformative prior to ensure the priors from our language models do not harm predictive performance.
By using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.
@@ 4 Limitations and future research
Within this work, we treat the language model as a domain expert to elicit prior distributions. These prior distributions need to be better than an uninformative prior to improve model performance and reduce data collection costs. A natural next step therefore is to question how these prior distributions compare to those elicited from domain experts, and whether language models provide an additional opportunity to reduce costs by supplementing expert elicited priors or by improving the expert elicitation framework. We began to explore how experts could be asked to produce more detailed task descriptions (Section 3.2 and Appendix A.10), but further settings would be interesting to explore.
As priors elicited from language models are a product of their training, they could contain unwanted bias. An approach to avoid this, would be to assign all features that could be considered sensitive to unwanted bias with an uninformative prior. However, prior distributions on other features might indirectly cause unwanted bias in a predictive model, and it is important to carefully think about whether these features should also be assigned an uninformative prior. Given an elicited prior distribution and observed examples, the Bayes factor can again be used for model selection to choose between combinations of an elicited prior with some or all features replaced with an uninformative prior. Additionally an expert, by providing information in the task description, could steer the language model away from unwanted bias in its provided distributions.
An interesting question is how we extend our results to neural networks. An immediate option is to ask a language model to suggest appropriate data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. A further exploration of methods would be beneficial.

### v2

@@ 2.3.3 Our contribution: Is the language model using the elicited prior in-context?
To quantify this, we use the energy statistic [33] (Appendix A.4), which measures how closely two distributions match (taking values $\in[0,1]$ with 00 indicating that two distributions are equal). Here, we test the difference between the parameter distributions from either the elicited distribution ($\theta\sim p_{M,T}(\theta)$) or the one approximated from the language model’s in-context predictive model ($\varphi\sim p_{M,T}(\varphi)$).
@@ 2.3.4 Our contribution: Inspecting the in-context posterior distribution
If the language model’s internal posterior is significantly different from the one calculated by applying Monte Carlo methods to the approximated prior, then in-context learning might be considered unusable for a given task, as it is not clear that the model is updating its prior belief about a predictive task from the data. This would suggest that the language model should not be used to generate posterior distributions on predictions in-context as it is improperly applying Bayes rule after observing demonstrations.
@@ 2.4 Bayes factor for model selection
To decide whether to use in-context learning or prior elicitation, given a data set, we consider the Bayes factor for model selection.
The Bayes factor compares the marginal likelihood ratio of two statistical models to determine under which model the data was most likely generated. Formally, given two models $\alpha_{0}$ and $\alpha_{1}$ with parameterisations $\theta_{0}$ and $\theta_{1}$, the Bayes factor calculated over a dataset $\mathcal{D}$ is:
@@
In our case, $\alpha_{0}$ refers to prior elicitation, with the prior determining the parameterisation $\theta$ of a predictive model, and $\alpha_{1}$ referring to in-context learning. We can approximate the Bayes factor using the means from samples of the likelihood of $D$ (Section 3.6).
We chose the Bayes factor over the Bayesian information criterion (BIC) because BIC uses the number of model parameters as part of the model selection. When comparing prior elicitation with in-context learning, BIC would almost always favour prior elicitation because the number of parameters required to make a prediction using in-context learning is of the order $10^{9}$.
@@ 2.5 Experimental datasets
Within our experiments, we present results on five real-world datasets and a synthetic dataset, covering both classification and regression tasks from various domains.
@@
To demonstrate the effectiveness of the LLM elicited priors on new tasks, we use a private dataset (i.e., not seen by the LLM) collected as part of a study on Dementia care which contains $10$ features of in-home activity and physiology and labels corresponding to days with a clinically validated labelling of positive or negative urinary tract infection (UTI), calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of data containing $78$ people and $371$ positive or negative days of UTI. This dataset has not appeared online and will not have been part of the language model’s training.
We then test four public benchmark datasets. The Breast Cancer dataset [36], where the task is to predict the diagnosis of a breast mass; the California Housing dataset [37], with the goal of predicting the median house value from US census data for a region; the Wine Quality dataset [38, 39] where the task is to predict the wine quality from its physicochemical characteristics; and the Heart Disease dataset [40, 41] where the task is to predict a heart disease diagnosis from physiological features.
@@ 3 Results
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo, with further language models tested in Appendices A.13.1 and A.13.2.
@@ 3.1 Evaluating the LLM-elicited priors
On the synthetic data, we learn to model $y=2x_{1}-x_{2}+x_{3}$ using a prior elicited from the language model. In this case, since we are testing how well the language model can provide priors based on informative task descriptions, we describe the exact relationships between $x$ and $y$ in the task descriptions. The LLM-elicited prior provides a lower MSE, by an order of magnitude, compared to an uninformative prior on the test set after $5$ observed examples and an equivalent MSE after $20$ data points are seen. This is likely because the task was simple enough to be learnt effectively by an uninformative prior after a few examples. However, it demonstrates that the language model is capable of providing effective prior distributions given a clear task description. In fact, the mixture of priors that the language suggests are shown in Appendix A.6.1 and are centred on the true parameter values, suggesting that the language model is able to convert its combined knowledge of a problem and a task description into useful prior distributions. The relationship between the details of the descriptions and the performance of elicited priors is studied within Appendix A.10.
For real-world datasets, we often see the same posterior performance improvements. On the Breast Cancer dataset, the improvement in accuracy is clear up to $40$ observed data points, where the mean accuracy is greater, and its variation is smaller. On the other hand, for the Wine Quality dataset, Figure 2(b) shows that the prior distribution on $\theta$ elicited from the language model is too strict, as it eventually reduces the mean accuracy of the logistic regression model when the number of training points is larger than $25$, whilst the variation in accuracy remains smaller. Given these samples, the likelihood of the prior distributions could have been measured (as described in 2.4) before the elicited distribution was used for new predictions.
@@
For both the California Housing and UTI datasets, the elicited prior provides a greater mean accuracy over the uninformative prior for all of the number of observed data points tested, illustrating the potential of using a language model for prior elicitation. This should be of particular note considering that the UTI dataset is private and will have never appeared within the language model’s training.
In general, we see that the greatest improvement in accuracy occurs for smaller numbers of observed samples, suggesting that the use of prior knowledge from language models is most relevant when users have access to only a few samples.
This can be important in scenarios where collecting labels is costly and time-intensive, as is the case in the UTI dataset. In Appendix A.6.2, we consider the accuracy achieved using uninformative priors and elicited priors in the context of collection time. We show here that a UTI predictive model might have achieved greater accuracy earlier in the study. In a clinical environment, this could have enabled improved UTI detection earlier, accelerating model development and validation.
In Appendix A.6, we present additional accuracy distributions (similar to Figure 2(a)), using the Heart Disease and Wine Quality datasets, which supplement the results in Figure 2(b). We further test the number of task descriptions required in Appendix A.11, and the quality of those descriptions in A.10.
Given two sets of priors, one elicited from a language model and an uninformative prior, the prior predictive distribution can be sampled and the likelihoods calculated given some data points. This can be done in a way similar to that described in Section 3.6.
@@ 3.2 Incorporating expert information
To study this, we chose to also include our own knowledge of the risk factors for UTI in the task description for prior elicitation [43]. In particular, we provide further information on three features available in the dataset that are positively correlated with the outcome. We added the following phrasing to the system prompts, based on clinical knowledge of UTIs: “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
@@
However, it is interesting that without this expert information, the LLM had already correctly identified the relationship between the risk of UTI and the frequency of the bathroom, nighttime disturbances, and the number of previous UTIs, as these are all positive predictors of UTI. The language model had already translated this information from its internal prior into the elicited distributions. This could explain the small improvement in posterior accuracy over LLM-elicited priors without additional expert information, shown in the line graph of Figure 3.
@@ 3.3 Extracting the in-context prior and posterior
The prior and posterior distributions estimated through a repeated sampling of the in-context predictions (as outlined in Section 2.3.2) are shown in Figures 4 and 5. Once we estimate the parameters of the linear model, we can use them to calculate the logits (or regression values) of the predictions returned by the language model (from in-context learning) and measure the error between the outputs of the language model and the estimated internal model. This is shown in the bottom row of Figures 4 and 5, suggesting that our estimates of the previous internal parameters are reliable for the real-world datasets. The upper row presents the distributions of the estimated parameters, with each colour representing a single feature.
For the prior distributions in Figure 4, we see variations in the estimated prior distributions across the datasets, suggesting that the language model uses some knowledge about the tasks when making predictions, even without any demonstration examples. However, we see that the language model often uses a prior centred around 00, suggesting that it is not particularly informative.
The extracted posterior distributions are shown in Figure 5 and show some differences compared to the prior distributions. For California Housing, Heart Disease, and Wine Quality, our estimation error increased but remained low. However, we find that the Breast Cancer dataset produces a large error for our predictions of the logits (Figure 5), suggesting that our estimates of the internal model parameters are not reliable for this case. It is interesting that, on the whole, we were able to produce a better approximation of the language model’s internal prior distribution than the posterior distribution, and this is likely due to the limitations of the language model and in particular, its capacity to understand demonstrations and instructions in the same prompt.
@@ 3.4 Is the language model being truthful?
Table 1 illustrates the differences between the extracted in-context prior distribution (Section 2.3.2) and the LLM-elicited distribution (Section 2.2). No dataset has similar distributions according to their energy distance, demonstrating that the language model uses a different prior distribution to make its own predictions compared to the one it provides for us to construct a separate predictive model.
These conclusions are supported in Appendix A.7, where we present histograms of the parameter values for each feature, which show that the two distributions visually differ significantly. This suggests that either the language model is unable to provide us with a reliable representation of its own prior knowledge or that it is unable to reliably use its own prior knowledge when making predictions. Since the in-context prior distributions are often centred around 00, we suspect that it is the latter and that the language model does not effectively use its own prior beliefs about a task in-context.
Since these priors disagree, we will investigate them further in Section 3.5, where we explore empirically whether in-context learning is approximating Bayesian inference and in Section 3.6, where we study decision criteria to understand which of these distributions is more likely to have generated the observed data using the Bayes factor.
@@ 3.5 Is in-context learning approximating Bayesian inference?
It is important that the model performs Bayesian inference, so that newly observed data points and prior knowledge are combined to form a posterior predictive distribution. In this section, we empirically evaluate the extent to which the language model is performing Bayesian inference in-context, by measuring the difference between the extracted in-context posterior distribution and Monte Carlo sampling applied to the extracted prior in-context distribution (Section 2.3.4).
@@
However, for Wine Quality and Heart Disease, the difference between the extracted posterior and the posterior samples of the extracted prior matches closely, and the error in our estimations of the parameters is low (Figures 4 and 5). In these cases, we are more confident that in-context learning is imitating Bayesian inference, and so should use a model selection tool to decide between prior elicitation and in-context learning.
Appendix A.8 presents histograms of the parameter values of each feature after Monte Carlo sampling or in-context learning have been performed. Figure 12 within Appendix A.8 supports the findings in this section and allows us to visualise the differences illustrated by Table 2.
These results provide empirical evidence that in-context learning sometimes approximates Bayesian inference, but not necessarily for all tasks.
@@ 3.6 Bayes factor and model selection
For in-context learning, we do not have access to $\varphi$ of the internal predictive model. However, since the language model provides us with predictions on targets given features directly, we can still approximate $p(\mathcal{D}|\alpha_{1})$ by repeated sampling of predictions. Here, we prompt the language model with our $100$ task descriptions and ask it to predict on $25$ data points from our dataset. We repeat this with $5$ different splits of the dataset to calculate the mean and standard deviation of the Bayes factor.
@@
For regression tasks, prior elicitation is preferred for $y=2x_{1}-x_{2}+x_{3}$, whilst for California Housing, in-context learning produces greater log-likelihoods. However, here, both in-context learning and prior elicitation provide a high prior mean squared error and low marginal likelihood, making the Bayes factor unreliable to compute. Furthermore, in the case of California Housing, in-context learning is not reliably estimating Bayesian inference (Section 3.5), suggesting that prior elicitation may still be preferable.
Whether prior elicitation or in-context learning provides prior distributions with greater likelihoods depends on the task, with both methods being the better choice over a uninformative prior for all cases tested. However, when factoring in the methods’ abilities to estimate Bayesian inference and the cost of producing predictions, prior elicitation is likely the better option for the cases tested.
@@ 3.7 Choosing between priors
As seen in Appendices A.13.1 and A.13.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this might not be the case for all datasets or language models. Given prior distributions elicited from different language models, we would like to choose the best one for our predictive task. For this, given some collected data, we can use the Bayes factor (Section 2.4) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques, such as the Bayesian information criterion. This is because we are performing the Bayesian inference ourselves, given the prior distributions. Furthermore, this method also allows us to compare the elicited distributions to the uninformative prior to ensure that the priors from our language models do not harm predictive performance. By using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.
@@ 4 Limitations and future research
In this work, we treat the language model as a domain expert to elicit prior distributions. These prior distributions need to be better than an uninformative prior to improve model performance and reduce data collection costs. A natural next step therefore is to question how these prior distributions compare to those elicited from domain experts, and whether language models provide an additional opportunity to reduce costs by supplementing expert elicited priors or by improving the expert elicitation framework. We began to explore how experts could be asked to produce more detailed task descriptions (Section 3.2 and Appendix A.10), but further tasks would be interesting to explore.
As priors elicited from language models are a product of their training, they could contain unwanted bias. An approach to avoid this would be to assign all features that could be considered sensitive to unwanted bias with an uninformative prior. However, prior distributions on other features might indirectly cause unwanted bias in a predictive model, and it is important to carefully think about whether these features should also be assigned an uninformative prior. Given an elicited prior distribution and observed examples, the Bayes factor can again be used for model selection to choose between combinations of an elicited prior with some or all features replaced with an uninformative prior. Additionally, by providing information in the task description, an expert could steer the language model away from unwanted bias in its provided distributions.
An interesting question is how we extend our results to neural networks. An immediate option is to ask a language model to suggest appropriate data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Further exploration of the methods would be beneficial.